OpenStudy (bruno102):

Write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2, -2/3)

OpenStudy (luigi0210):

Work backwards, you already have the solutions: x=0 x=1

OpenStudy (luigi0210):

You wouldn't happen to know derivatives by any chance? @bruno102

OpenStudy (bruno102):

no sorry that's the question

OpenStudy (bruno102):

This question is really confusing me..

OpenStudy (anonymous):

y=x(x-1)=x^2-x

OpenStudy (anonymous):